# Resource Public Key Infrastructure BGP Border Gateway Protocol

### v1

@@ Abstract
Resource Public Key Infrastructure (RPKI) is a critical security mechanism for BGP, but the complexity of its architecture is a growing concern as its adoption scales. Current RPKI design heavily reuses legacy PKI components, such as X.509 EE-certificates, ASN.1 encoding, and XML-based repository protocols, all these introduce excessive cryptographic validation, redundant metadata, and inefficiencies in both storage and processing. We show that these design choices, although based on established standards, create significant performance bottlenecks, increase the vulnerability surface, and hinder scalability for wide-scale Internet deployment.
@@ I Introduction
According to NIST monitor [^1] https://rpki-monitor.antd.nist.gov as of April 2025, more than 55% of the Internet prefixes are certified with RPKI and according to Rovista [^2] https://rovista.netsecurelab.org/analytics almost 30% of networks validate BGP announcements. With the growing adoption of RPKI, it also plays an increasingly central role to the stability and resilience of the Internet. Recently, RPKI also gained political importance: according to the routing security roadmap published by the US White House in 09/2024, RPKI is critical for national security [8]. The roadmap identifies RPKI as a mature, ready-to-implement technology to mitigate vulnerabilities in BGP, and recommends deployment on all networks. It is thus expected RPKI deployment will continue to grow.
@@
RPKI was standardized relatively late compared to other protocols, like HTTPS (RFC9110) or DNSSEC (RFC9364), allowing authors of the RPKI specification to learn from issues in other technologies and incorporate existing solutions into the design of RPKI. Observing that building a new technology from scratch is labor-intensive and error-prone, RPKI utilizes many existing concepts from other protocols (RFC6480), like a Public Key Infrastructure (PKI), (RFC5280), X.509 certificate templates and X.680 DER-encoding used in TLS (RFC8446), and a distributed delegated server architecture similar to DNS (RFC1035). Designing RPKI on-top established technologies allows implementations to use existing tooling, like OpenSSL for parsing and validation, facilitating faster development and increasing acceptance by the community.
@@
However, RPKI is not merely a replica of other technologies; it aims to address the shortcomings of the fundaments it builds upon. For example, certificate revocation in TLS has historically proved ineffective with many TLS implementations not checking for revoked certificates at all due to the large client-side overhead [10]. RPKI overcomes this issue by tightly integrating the Certificate Revocation List (CRL) in its architecture and mandating all clients to check it before validation (RFC6487). RPKI also acknowledges the issue of diverging certificate structure over time, resulting from the high flexibility for certificate issuers in TLS, and instead introduces multiple tight requirements to certificate structure (RFC6487). While these design decisions improve security, they make RPKI substantially complex, and make validation of certificates in RPKI more computationally intensive.
@@
In this paper we investigate the architectural complexity of RPKI and analyze how its foundational design choices, such as reuse of legacy PKI structures, redundant cryptographic layers, and verbose data encodings, create systemic hurdles and challenges for RPKI performance, especially during validation, for correctness of the implementation, resilience of RPKI deployments and future scalability. In addition, the complexity of the specification makes RPKI challenging to implement, leading to vulnerabilities [21, 13] and reducing the diversity of available RPKI implementations (only 1 open-source repository software and 3 RP implementations are widely deployed). Contrary to previous work, rather than attributing issues in RPKI to implementation flaws or insufficient resources [21, 14], we show that many of the issues in RPKI stem from inherited complexity embedded in its object formats, certificate handling, and repository protocols. We perform a comprehensive analysis of RPKI relying party (RP) validator, correlating processing overhead with specific features mandated by the RFCs, and explore how the respective design choices result in intensive processing requirements for the software. We reveal that over half of validation time is consumed by non-essential certificate parsing and signature verification. Our evaluations on empirically derived datasets show that the complexity and inefficiency will further exacerbate as more RPKI objects are issued, more RPKI validators are adopted and more use-cases, like ASPA, are incorporated, eventually leading to extremely resource-intensive and instable deployments and operation.
@@
Based on these findings, we derive major improvements to the RPKI specification that reduce complexity, bandwidth requirements, object sizes and computational effort for implementations: We propose iRPKI, a redesign of RPKI that preserves the security and trust guarantees of the existing system while substantially simplifying its architecture. iRPKI introduces compact object formats, removes the need for per-object EE-certificates and redundant signatures, merges revocation and integrity mechanisms, and replaces inefficient encodings like XML and ASN.1 with lightweight alternatives such as Protobuf. We implement our improvements to the design of RPKI into a Routinator software package and experimentally show that our suggested changes not only reduce processing time by up to 93%, and similarly bandwidth overhead by 95%, cut the RRDP file sizes by 95%, shrink cache memory usage by 88%, but also eliminate entire classes of vulnerabilities rooted in parsing and validation logic, e.g., those related to EE-certificate parsing and redundant signature checks, that account for the majority of known RPKI implementation flaws [13]. We open-source iRPKI objects, the code of iRPKI based on the adapted Routinator code, and the repository tool [^3] Study Data.
@@
Paper structure. We review related work in Section II. We review RPKI in Section III and analyze its architecture, mapping the complexities in the current RPKI design and implementations, mapping them to RFCs, in Section V. We analyze the effects of the complexity factors in the real world in Section IV and illustrate how iRPKI overcomes them in Section VI. We evaluate iRPKI in VII, discuss implications of our work in VIII. We conclude in Section IX.
@@ II Related Work
Previous research on RPKI explored vulnerabilities in RPKI implementations and errors in operation and proposed countermeasures [16, 18, 6, 14]. We next compare our work to previous research that developed mitigations and to papers that explored problems in RPKI, that we in this work trace to complexity in RPKI design. [6] conducted an analysis of thresholds in RPKI software. Their work showed that the growing size of RPKI repositories lead to operational problems for RPs, but did not identify that the underlying cause for the large RPKI repository sizes were inefficiencies in the RPKI design. Our work proposes to fix the underlying factors that contribute to the growth of repositories beyond manageable limits, by improving the RPKI design and reducing the size of repositories by a factor of 18x, which enables smaller thresholds, as fetch times are decreased.
[13] developed a fuzzing tool for RPKI software, and detected 18 vulnerabilities in RPs. The authors attributed the factors to lack of test tools for RPKI software developers, but did not identify the issues in the fundamental complexity of RPKI architecture. Our proposed design with iRPKI reduces the complexity of objects and RRDP, improving efficiency and reducing the susceptibility to bugs in parsing the complex RPKI objects. As we discuss in Section V, 10 of the vulnerabilities found by [13] would have been eliminated if iRPKI were used, since the improved object structures remove the fields and structure that triggered the vulnerability.
@@ III Overview of RPKI
Ownership. The design of RPKI follows the design of existing PKIs: Binding ownership of resources to a cryptographic key. In the context of RPKI, the ownership of IP prefixes and Autonomous System Numbers (ASNs) is bound to the ownership of a key through a signed certificate - similar to how HTTPS binds ownership of domains to a key through a signed certificate. To validate ownership of resources, when certificates are issued and signed by a Certificate Authority (CA), RPKI follows the Internet number resource allocation hierarchy defined by IANA (top level), then Regional Internet Registries (RIRs) who allocate resources to entities like Internet Service Providers (ISPs), and so on - this hierarchy matches the way how resources are actually assigned in the Internet. The design of RPKI follows this delegation tree to establish trust: IANA anchors the global legitimacy of network resources delegation in RPKI. The five RIRs act as trust-roots and sign certificates for allocated resources. Owners of resources can further sub-allocate their resources, and use their key to sign the child certificate to attest this allocation. In RPKI, each certificate owner is a CA and can sign child certificate. Other entities can validate the ownership by following a trust chain, starting at the CA and validating each parent signature until validation reaches a trust root.
@@
Validation. RPKI utilizes a middleware called RPs to fetch RPKI data from repositories. RPs access the global RPKI data, validate all objects, and compile a list of the validated data, called Validated ROA Payloads (VRPs). Downloading data from repositories is implemented through two access protocols. Initially, rsync was used to fetch data from remote repositories. Eventually, a second transport mechanism, RPKI Repository Delta Protocol (RRDP, RFC8182), was standardized. RRDP allows incremental fetches of repository data by including incremental changes in Delta.xml files. The RP software is informed over the current repository state through a serial number and session id inside a Notification.xml file that is linked in the certificate of all CAs publishing at a given repository. If an update over deltas is possible, the deltas are applied to the local state incrementally until the current serial number is reached. If the delta update fails or no local state is available, e.g., in the first start-up of an RP, the entire repository state is downloaded through a Snapshot.xml file. The snapshot contains the names and base64-encoded binary content of all objects in the repository. Routers download the VRPs from RPs through the RPKI to Router protocol (RTR) and use it to validate incoming BGP messages.
@@ IV-A Analysis Methodology
To ensure reproducibility of our results and reduce the outside noise of our evaluations, we create a local, isolated setup, and run our analysis in it. This is important, as real-world processing times are heavily influenced by external factors, like configuration errors or networking issues in remote repositories [14]. For our setup, we create a local RPKI repository with 100 test CAs. We choose 100, since we experimentally find that the flamegraph results do not change when including more CAs. Each CA includes one manifest and one CRL. For each CA, we additionally include 6 ROAs, the average amount of ROAs per CA in real-world RPKI. To serve our local repository, we use an nginx webserver without bandwidth limitations. Since download times relate to object size, looking at the size of objects allows us to quantify impact on download times, omitting the need to limit download speed in our flamegraph analysis. We explain how our results generalize to real-world RPKI deployments, including limited download speed and large amount of CAs in Section VII.
Flamegraph results. The results of our evaluation with Routinator RP are shown in Figure 1. The majority of processing time in the RP is spent to validate public key signatures of the signed objects totaling to 62% of RP processing. This processing time is spent to validate 1200 signature on ROAs (2 per object), 200 signatures on manifests (2 per objects), and 100 signatures on CRLs and CA certificates (1 per object). Both ROAs and manifests contain 2 signatures per object since they contain a certificate and a signed object signature. For example, looking at ROAs (47% of total processing time) Routinator spends 24% of total processing time on validating the certificate signature and 23% to validate the ROA content signature. However, not all processing time is spend on public key validation. 13% of processing time is spent on parsing and validating non-signature fields of signed objects (ROAs and MFTs). Further, parsing and validating the RRDP XML files takes 9%, with the large majority of the time spent on parsing the Snapshot. HTTPS request handling, like socket creation, takes about 4% of total processing time. Finally, the RP takes 12% of processing time in other functions, like internal state management, thread management, and disc writes. In conclusion, the RP spends the large majority of time on cryptographic validation, followed by parsing of objects and snapshot files.
@@ IV-B Complexity of EE-Certificates
This design is supported by the ASN.1 type signedObject (RFC5652), allowing to add a signature on encapsulated content, and optionally include signer certificates. Adding the signing certificate directly to objects is mainly motivated by self-sufficiency: According to RFC5652 Section 5.1, this design allows the object to contain all certificates necessary to validate it up to a trust-root, making it fully self-sufficient in validation. The RFC acknowledges that adding the full chain is not required, and objects may contain only a subset of certificates since other necessary certificates for path validation can also be attained through other means. The latter option is used in RPKI: RPKI objects include a single certificate that is only used to validate this specific object content. All other certificates for path validation up to the trust root are obtained through the RPKI architecture and are not contained within the signed object.
@@
This hybrid design (only containing a subset of required certificates) raises the question why RPKI includes any certificates in objects instead of using the other validation option given by RFC5652, which is omitting any contained certificates and having the content directly signed by the CA.
@@
While not explicitly stated as the reason in the RFC, this design is likely based on supporting potential future use-cases. Including EE-certificates in each object allows more independent distribution of signed objects, as it makes objects partially self-sufficient: The RP can use certificate fields for content validation, and to find issuer certificates for path validation. This potential use-case is specifically mentioned in RFC6480 §3.1, describing ROAs might be distributed e.g., within BGP messages. However, to-date, out-of-band distribution of ROAs is neither supported by the specification, nor is the creation or processing of such objects supported by any current repository or RP implementation. To the best of our knowledge, no current efforts within the IETF or software development community are working on such a use-case.
@@
Including the EE-certificate also provides benefits in the current architecture. RFC6487 states that EE-certificates allow targeted revocation of a signed object through its certificate. By using EE-certificates, revocation of objects can use the serial number of the certificate within the signed object instead of having to revoke the CA certificate, re-issuing it and re-signing all child objects. It improves revocation efficiency. Using EE-certificates has the additional advantage of re-usability of existing technologies and open-source solutions (RFC6488). For example, attesting additional aspects of an object (like issuer name) is implemented through existing certificate extensions instead of developing new concepts for storage within object content. By using EE-certificates, RPKI thus aims to lower the complexity of implementing new object templates by using existing (complex) template definitions where possible.

### v2

@@ Abstract
Resource Public Key Infrastructure (RPKI) is a critical security mechanism for BGP, but the complexity of its architecture is a growing concern as its adoption scales. Current RPKI design heavily reuses legacy PKI components, such as X.509 EE-certificates, ASN.1 encoding, and XML-based repository protocols, which introduce excessive cryptographic validation, redundant metadata, and inefficiencies in both storage and processing. We show that these design choices, although based on established standards, create significant performance bottlenecks, increase the vulnerability surface, and hinder scalability for wide-scale Internet deployment.
@@ I Introduction
According to NIST monitor [30] as of April 2025, more than 55% of the Internet prefixes are certified with RPKI and according to Rovista [21] almost 30% of networks validate BGP announcements. With the growing adoption of RPKI, it also plays an increasingly central role to the stability and resilience of the Internet. Recently, RPKI also gained political importance: according to the routing security roadmap published by the US White House in 09/2024, RPKI is critical for national security [14]. The roadmap identifies RPKI as a mature, ready-to-implement technology to mitigate vulnerabilities in BGP, and recommends deployment on all networks. It is thus expected RPKI deployment will continue to grow.
@@
RPKI was standardized relatively late compared to other protocols, like HTTPS [32] or DNSSEC [13], allowing authors of the RPKI specification to learn from issues in other technologies and incorporate existing solutions into RPKI design. Observing that building a new technology from scratch is labor-intensive and error-prone, RPKI utilizes many existing concepts from other protocols [20], like a Public Key Infrastructure [6], X.509 certificate templates and X.680 DER-encoding [32], and a distributed delegated server architecture similar to DNS [28]. Designing RPKI on-top established technologies allows implementations to use existing tooling, like OpenSSL for parsing and validation, facilitating faster development and increasing acceptance by the community.
@@
However, RPKI is not merely a replica of other technologies; it aims to address the shortcomings of the fundaments it builds upon. For example, certificate revocation in TLS has historically proved ineffective with many TLS implementations not checking for revoked certificates at all due to the large client-side overhead [22]. RPKI overcomes this issue by tightly integrating the Certificate Revocation List (CRL) in its architecture and mandating all clients to check it before validation [18]. RPKI also acknowledges the issue of diverging certificate structure over time, resulting from the high flexibility for certificate issuers in TLS, and instead introduces multiple tight requirements to certificate structure [18]. While these design decisions improve security, they make RPKI substantially complex, and make validation of certificates in RPKI more computationally intensive.
@@
In this paper we investigate the architectural complexity of RPKI and analyze how its foundational design choices, such as reuse of legacy PKI structures, redundant cryptographic layers, and verbose data encodings, create systemic hurdles and challenges for RPKI performance. We identify these issues across RPKI components, like validation complexity, correctness of implementations, resilience of RPKI deployments and future scalability. In addition, the complexity of the specification makes RPKI challenging to implement, leading to vulnerabilities [40, 25] and reducing the diversity of available RPKI implementations (only 1 open-source repository software and 3 RP implementations are widely deployed). Contrary to previous work, rather than attributing issues in RPKI to implementation flaws or insufficient resources [40, 26], we show that many of the issues in RPKI stem from inherited complexity embedded in its object formats, certificate handling, and repository protocols. We perform a comprehensive analysis of RPKI Relying Party (RP) validators, correlating processing overhead with specific features mandated by the RFCs, and explore how the respective design choices result in intensive processing requirements for the software. We reveal that over half of validation time is consumed by non-essential certificate parsing and signature verification. Our evaluations on empirically derived datasets show that the complexity and inefficiency will further exacerbate as more RPKI objects are issued, more RPKI RPs are adopted and more use-cases, like ASPA, are incorporated, eventually leading to extremely resource-intensive and instable deployments and operation.
@@
Based on these findings, we derive major improvements to the RPKI specification that reduce complexity, bandwidth requirements, object sizes and computational effort for implementations: We propose iRPKI, a redesign of RPKI that preserves the security and trust guarantees of the existing system while substantially simplifying its architecture. iRPKI introduces compact object formats, removes the need for per-object EE-certificates and redundant signatures, merges revocation and integrity mechanisms, and replaces inefficient encodings like XML and ASN.1 with lightweight alternatives such as Protobuf. We implement our improvements to RPKI design into the Routinator software package and experimentally show that our suggested changes not only reduce processing time by up to 93%, and similarly bandwidth overhead by 95%, cut the RRDP file sizes by 95%, shrink cache memory usage by 88%, but also eliminate entire classes of vulnerabilities rooted in parsing and validation logic, e.g., those related to EE-certificate parsing and redundant signature checks, that account for the majority of known RPKI implementation flaws [25]. We open-source iRPKI objects, the code of iRPKI based on the adapted Routinator code, and the repository tool111www.dropbox.com/scl/fi/pvs6odh2ml25azifd6vg9/submission_ndss.zip.
@@
Paper structure. We review related work in Section II. We discuss RPKI in Section III and analyze its architecture, mapping the complexities in the current RPKI design and implementations, mapping them to RFCs, in Section V. We analyze the effects of the complexity factors in the real world in Section IV and illustrate how iRPKI overcomes them in Section VI. We evaluate iRPKI in VII and discuss implications of our work in VIII. We conclude in Section IX.
@@ II Related Work
Previous research on RPKI explored vulnerabilities in RPKI implementations and errors in operation and proposed countermeasures [29, 37, 11, 26]. We next compare our work to previous research that developed mitigations and to papers that explored problems in RPKI, that we in this work trace to complexity in RPKI design. [11] conducted an analysis of thresholds in RPKI software. Their work showed that the growing size of RPKI repositories leads to operational problems for RPs, but did not identify that the underlying cause for the large RPKI repository sizes were inefficiencies in the RPKI design. Our work proposes to fix the underlying factors that contribute to the growth of repositories beyond manageable limits, by improving the RPKI design and reducing the size of repositories by a factor of 18x, which enables smaller thresholds, as fetch times are decreased.
[25] developed a fuzzing tool for RPKI software, and detected 18 vulnerabilities in RPs. The authors attributed the vulnerabilities to a lack of test tools for RPKI software developers, but did not identify the issues in the fundamental complexity of RPKI architecture. Our proposed design with iRPKI reduces the complexity of objects and RRDP, improving efficiency and reducing the susceptibility to bugs in parsing the complex RPKI objects. As we discuss in Section V, 10 of the vulnerabilities found by [25] would have been eliminated if iRPKI were used, since the improved object structures remove the fields and structure that triggered the vulnerability.
@@ III Overview of RPKI
Ownership. The design of RPKI follows the design of existing PKIs: Binding ownership of resources to a cryptographic key. In the context of RPKI, the ownership of IP prefixes and Autonomous System Numbers (ASNs) is bound to the ownership of a key through a signed certificate - similar to how HTTPS binds ownership of domains to a key through a signed TLS certificate. To validate ownership of resources, when certificates are issued and signed by a Certificate Authority (CA), RPKI follows the Internet number resource allocation hierarchy defined by IANA (top level), then Regional Internet Registries (RIRs) who allocate resources to entities like Internet Service Providers (ISPs), and so on - this hierarchy matches the way how resources are actually assigned in the Internet. The design of RPKI follows this delegation tree to establish trust: IANA anchors the global legitimacy of network resources delegation in RPKI. The five RIRs act as trust-roots and sign certificates for allocated resources. Owners of resources can further sub-allocate their resources, and use their key to sign the child certificate to attest this allocation. In RPKI, each certificate owner is a CA and can sign child certificates. Other entities can validate the ownership by following a trust chain, starting at the CA and validating each parent signature until validation reaches a trust root.
@@
Validation. RPKI utilizes a middleware called RPs to fetch RPKI data from repositories. RPs access the global RPKI data, validate all objects, and compile a list of the validated data, called Validated ROA Payloads (VRPs). Downloading data from repositories is implemented through two access protocols. Initially, rsync was used to fetch data from remote repositories. Eventually, a second transport mechanism, RPKI Repository Delta Protocol (RRDP), was standardized [4]. RRDP allows incremental fetches of repository data by including incremental changes in Delta.xml files. The RP software is informed of the current repository state through a serial number and session id inside a Notification.xml file that is linked in the certificate of all CAs publishing at a given repository. If an update over deltas is possible, the deltas are applied to the local state incrementally until the current serial number is reached. If the delta update fails or no local state is available, e.g., in the first start-up of an RP, the entire repository state is downloaded through a Snapshot.xml file. The snapshot contains the names and base64-encoded binary content of all objects in the repository. Routers download the VRPs from RPs through the RPKI to Router protocol (RTR) and use it to validate incoming BGP messages.
@@ IV-A Analysis Methodology
To ensure reproducibility of our results and reduce the outside noise of our evaluations, we create a local, isolated setup, and run our analysis in it. This is important, as real-world processing times are heavily influenced by external factors, like configuration errors or networking issues in remote repositories [26]. For our setup, we create a local RPKI repository with 100 test CAs. We choose 100, since we experimentally find that the flamegraph results do not change when including more CAs. Each CA includes one manifest and one CRL. For each CA, we additionally include 6 ROAs, the average amount of ROAs per CA in real-world RPKI. To serve our local repository, we use an nginx webserver without bandwidth limitations. Since download times relate to object size, looking at the size of objects allows us to quantify impact on download times, omitting the need to limit download speed for analysis. We explain how our results generalize to real-world RPKI deployments, including limited download speed and large amount of CAs in Section VII.
Flamegraph results. The results of our evaluation with Routinator RP are shown in Figure 1. The majority of processing time is spent to validate the signed objects, totaling to 62% of RP processing. This processing time is spent to validate 1200 signature on ROAs (2 per object), 200 signatures on manifests (2 per objects), and 100 signatures on CRLs and CA certificates (1 per object). Both ROAs and manifests contain 2 signatures per object since they contain a certificate and a signed object signature. For example, looking at ROAs (47% of total processing time) Routinator spends 24% of total processing time on validating the certificate signature and 23% to validate the ROA content signature. However, not all processing time is spend on public key validation. 13% of processing time is spent on parsing and validating non-signature fields of signed objects (ROAs and MFTs). Further, parsing and validating the RRDP XML files takes 9%, with the large majority of the time spent on parsing the Snapshot. HTTPS request handling, like socket creation, takes about 4% of total processing time. Finally, about 12% of processing time is spent in other functions, like internal state management, thread management, and disc writes. In conclusion, the RP spends the large majority of time on cryptographic validation, followed by parsing of objects and snapshot files.
@@ IV-B Complexity of EE-Certificates
This design is supported by the ASN.1 type signedObject [15], allowing to add a signature on encapsulated content, and optionally include signer certificates. Adding the signing certificate directly to objects is mainly motivated by self-sufficiency: According to [15] Section 5.1, this design allows the object to contain all certificates necessary to validate it up to a trust-root, making it fully self-sufficient in validation. The RFC acknowledges that adding the full chain is not required, and objects may contain only a subset of certificates since other necessary certificates for path validation can also be attained through other means. The latter option is used in RPKI: RPKI objects include a single certificate that is only used to validate this specific object content. All other certificates for path validation up to the trust root are obtained through the RPKI architecture and are not contained within the signed object.
@@
This hybrid design (only containing a subset of required certificates) raises the question why RPKI includes any certificates in objects instead of using the other validation option given by [15], which is omitting any contained certificates and having the content directly signed by the CA.
@@
While not explicitly stated as the reason in the RFC, this design is likely based on supporting potential future use-cases. Including EE-certificates in each object allows more independent distribution of signed objects, as it makes objects partially self-sufficient: The RP can use certificate fields for content validation, and to find issuer certificates for path validation. This potential use-case is specifically mentioned in [20] §3.1, describing ROAs might be distributed, e.g., within BGP messages. However, to-date, out-of-band distribution of ROAs is neither supported by the specification, nor is the creation or processing of such objects supported by any current repository or RP implementation. To the best of our knowledge, no current efforts within the IETF or software development community are working on such a use-case.
@@
Including the EE-certificate also provides benefits in the current architecture. [18] states that EE-certificates allow targeted revocation of a signed object through its certificate. By using EE-certificates, revocation of objects can use the serial number of the certificate within the signed object instead of having to revoke the CA certificate, re-issuing it and re-signing all child objects. It improves revocation efficiency. Using EE-certificates has the additional advantage of re-usability of existing technologies and open-source solutions [19]. For example, attesting additional aspects of an object (like issuer name) is implemented through existing certificate extensions instead of developing new concepts for storage within object content. By using EE-certificates, RPKI thus aims to lower the complexity of implementing new object templates by using existing (complex) template definitions where possible.